# A Low-Complexity Speech Codec Using Parametric Dithering for ASR

### v2

@@ 3 Analysis
Given the reliance of ASR models on feature extraction from log-mel spectrograms and prior work investigating the impact of spectral distortion on ASR performance [radford2022robust], we consider the structural integrity of Time-Frequency (TF) features “visible” to models as a baseline for optimal dither design. For an error signal $\varepsilon$ of total length NN with discrete time and frequency indices $n$ and $f$, we adopt two metrics to characterize this integrity: one quantifies the overall increase in quantization error power (MSE), and the other captures the spectral distribution of the error, both of which are represented by the Power Spectral Density (PSD) as $S_{\varepsilon\varepsilon}(f)=\left|\sum_{n=0}^{N}\varepsilon(n)e^{-j2\pi fn/N}\right|^{2}$. We aim for a smooth error power spectrum and minimal frequency energy increase, to best ensure retention of the critical TF patterns that ASR models are trained on.
@@ 4 Codec
The proposed codec follows the pipeline outlined in Fig. 4. Input and dither signals of NN samples are summed before being passed through the mid-tread $b$-bit quantizer. This is followed by Huffman Coding, an optimal lossless compression technique which achieves an average code length within 1 bit of the source entropy.

### v3

@@ 3 Analysis
Given the reliance of ASR models on feature extraction from log-mel spectrograms and prior work investigating the impact of spectral distortion on ASR performance [radford2022robust], we consider the structural integrity of Time-Frequency (TF) features “visible” to models as a baseline for optimal dither design. For an error signal $\varepsilon$ of total length $N$ with discrete time and frequency indices $n$ and $f$, we adopt two metrics to characterize this integrity: one quantifies the overall increase in quantization error power (MSE), and the other captures the spectral distribution of the error, both of which are represented by the Power Spectral Density (PSD) as $S_{\varepsilon\varepsilon}(f)=\left|\sum_{n=0}^{N}\varepsilon(n)e^{-j2\pi fn/N}\right|^{2}$. We aim for a smooth error power spectrum and minimal frequency energy increase, to best ensure retention of the critical TF patterns that ASR models are trained on.
@@ 4 Codec
The proposed codec follows the pipeline outlined in Fig. 4. Input and dither signals of $N$ samples are summed before being passed through the mid-tread $b$-bit quantizer. This is followed by Huffman Coding, an optimal lossless compression technique which achieves an average code length within 1 bit of the source entropy.